Homework/ Class Exercise on Topic 6 to 8 Goodyear Tires Malaysia which located at Subang Jaya currently produces tires at their plant during two 12-hour shifts. The night-shift employees are planning to ask for a raise because they believe they are producing more tires per shift than the day-shift. The production supervisor randomly selected some daily production runs from each shift with the results given below (in 1,000s of tires produced). a) Shift Production (in 1,000s) Day | 107.5 | 118.6 | 124.6 | 101.6 | 113.6 | 119.6 ļ 1206 Tì09. | 105,9 Night 115.6 1094 121.6 1287 136.6 125.4 121.3 180.6 17.5 At 5 percent significance level, do these data indicate that the night shift is producing more tires per shift?
Answer #1

H0: Null Hypothesis: mu 1geq mu 2

HA: Alternative Hypothesis: 111 < μ2

From the given data, the following statistics are calculated:

n1= 9

ar{x}1 = 113.51

s1 = 7.8058

n2 = 9

ar{x}2 = 128.52

s2 = 21.0397

SP nl n2- 2

(9-1) × 7.80582 + (9-1) × 21.03972 9+9- 2 = V25 1.8-15.8682

SE=s_{P}sqrt{1/n1+1/n2}

15.8682 V1/94 1 /9 7.4803

Test Statistic is:

t = (113.51 - 128.52)/7.4803 = - 2.0066

alpha = 0.05

ndf =9 +9 - 2 = 16

One Tail - Left SideTest

From Table, critical value of t = - 1.7459

Since the calculated value of t = - 2.0066 is less than critical value of t = - 1.7459, the difference is significant. Reject null hypothesis.

Conclusion:

The data support the claim that the night shift is producing more tires per shift.